OpenStudy (anonymous):

On a true-false test, the ratio of true answers to false answers is 4 to 3. There are 18 false answers. How many answers are true? A.2. B.14. C.20. D.24.

OpenStudy (hoblos):

T/F = 4/3 for F=18 T= F(4/3) = 24

OpenStudy (anonymous):

24 true

OpenStudy (anonymous):

The ratio is 4 to 3, which is written as 4/3. You can set this up as an equivalence: 4/3 = x/18